OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? - 5 over 3 5 over 3 -3 over 5 3 over 5 hellp I tried out the answer and i got the 3rd answer is that correct or no ?

OpenStudy (anonymous):

yes you are.

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (2,3) x1= 2 and y1 = 3 (-3,0) x2 = -3 and y2 = 0 now lets sub m = (0 - 3) / (-3 - 2) m = -3/-5 m = 3/5 you are not correct. The answer is D.

OpenStudy (anonymous):

Thank you for explaining it @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

:)